OpenStudy (morganwarren):

If f(x) = 2x + 3 and g(x) = x2 + 1,find f(g(3))

jhonyy9 (jhonyy9):

so first calculi g(3) = ?

OpenStudy (morganwarren):

I have no clue how to do this. I'm taking a pretest.

jhonyy9 (jhonyy9):

so just substitute the 3 in place of x inside g(x) = x^2 +1

OpenStudy (morganwarren):

So 3^2+1?

jhonyy9 (jhonyy9):

yes = ?

OpenStudy (morganwarren):

10

jhonyy9 (jhonyy9):

yes - so now you know that g(3) = 10 and you need to calculi f(g(3)) so what will be f(10) = ? - because g(3) = 10 - yes ?

OpenStudy (morganwarren):

would it be 2(3)+3?

jhonyy9 (jhonyy9):

so f(10) will get same g(3) just with difference that substitute in this case inside f(x) in place of x 10

OpenStudy (morganwarren):

23?

jhonyy9 (jhonyy9):

how you have got it ?

jhonyy9 (jhonyy9):

but yes is correct sure f(g(3)) = 2*10 +3=23

OpenStudy (morganwarren):

Okay thank you!

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (morganwarren):

Yes i do!

jhonyy9 (jhonyy9):

was my pleasure anytime good luck bye